OpenStudy (anonymous):

A colony of bacteria is growing exponentially, doubling in size every 50 minutes, How many minutes will it take for the colony to become six times it current size?

OpenStudy (anonymous):

Do you know the formula for doubling growth rate?

OpenStudy (anonymous):

the exponential growth and doubling equation is: P(t)=Po(2)^(t-to)/d

OpenStudy (anonymous):

\[6=2^{\frac{t}{50}}\] solve for \(t\)

OpenStudy (anonymous):

takes exactly two steps: change of base gives \[\frac{t}{50}=\frac{\ln(6)}{\ln(2)}\] \[t=\frac{50\ln(6)}{\ln(2)}\]

OpenStudy (anonymous):

what would the exact answer be and how would you solve that, i want to understand each step.